# What’s DAT? Three Case Studies of Measuring Software Development Productivity at Meta With Diff Authoring Time

### v1

@@ 7.3. Results
We initially applied this method to two major code-sharing frameworks, covering several thousand of code-sharing diffs and counterfactual diffs monthly. Our results indicate that in this case alone, we save thousands of hours of DAT annually, a relative improvement to non-shared development of over 50%. When we set out, it was unclear whether there were substantial savings from code sharing; DAT enabled a precise statistical comparison, moving beyond anecdotes and intuition.
@@ 8. Threats to Validity
As a top-line metric DAT strikes a balance between support of all engineering workflows at the company (external validity), and precision for each diff (internal validity). These objectives are often at odds, as the example of Anchor-DAT shows: extrapolation increases tool coverage at the expense of potentially covering activity not related to this diff.
To mitigate this, we performed a quadruple validation of DAT (see Section 4), including large-scale surveys and end-to-end tests with GT (construct validity). Moreover, we plan to increase accuracy by supporting more tools and IDEs in a native way. For instance, we have already seen the beneficial effects of the addition of Android Studio as a natively supported IDE.
@@ 9.1. Going beyond velocity
This paper demonstrates the broad applicability of DAT in practice, enabling Software Engineering-related case studies and experiments on different levels, from language features and compiler improvements to a generalized method to calculate framework efficiency gains. While DAT offers insights into velocity, there are other important aspects to software development such as quality (Cavano and McCall, 1978). Our experiments and case studies have shown, though, that these often go together: a DAT improvement frequently co-occurs with other desired properties, such as more static safety guarantees through typed mocking in Section 5).
Our studies have established that language and framework improvements (e.g., the React compiler improvements in Section 6) provide large velocity improvements. However, for an end-to-end analysis, their initial development efforts need to be balanced against the later DAT savings, ensuring the benefits outweigh the initial added development overhead. Fortunately the initial development cost can be measured with DAT, too, providing a net benefit analysis of these features. Having performed over 20 DAT experiments and case studies, we are now in a position to inform planning and prioritization of development in a more rigorous, data-driven way.
9.2. Dealing with small effect sizes
Another important consideration is the challenge of measuring a statistically significant change in DAT, especially when working with small sample sizes. To overcome this, experiments often need to run for extended periods to account for variability in workflows and usage patterns. This distinguishes internal developer experimentation from external consumer experimentation, where changing the position of a button by a few pixels can yield measurably different results (Goodui.org, 2020). The fidelity of DAT might not lend itself to such types of features or improvements.
We believe that this could also be attributed to the fact that developers are highly cognitively engaged in their work and constantly striving to find optimizations themselves. Their optimization scope and creativity are much larger than the often confined ecosystems in the consumer-facing product world, which equally means that capturing them in a metric makes it so much harder. An example of this would be developers that continue to use their favorite bespoke editor, even if there is only one officially supported workflow to creating diffs, and copy their edits in for example. The DAT for such copy-and-paste might be extremely short, and thus be of no value for experimentation.
In such cases, engineering teams might show impact by improving an operational metric more dependent on their changes, for example, number of successful searches. While a statistically significant, positive change in this metric might not yield statistically significant effects on DAT, the team could rely on a previously demonstrated, principal correlation between that metric and DAT to claim productivity improvements.
9.3. Embracing an experiment-first culture
An experimentation workflow requires time and rigor, and may even require a shift in the behavior of Infrastructure engineers, who might not be used to this mindset. As an example, the experiment in Section 5 took into account the DAT experiment and involved a staged migration of mock call-sites, to ensure DAT would be correctly measured. This additional effort is often only worth the overhead if engineers expect a significant DAT reduction. This however assumes that a feature is known to be strictly beneficial a-priori, which is often only a post-hoc learning. To design code in a DAT-first way can significantly alter the feature itself, as was the case in Section 5, where we did not migrate all untyped mock call sites at once. In this sense, there could be an opportunity cost to running an experiment.
Finally, DAT can serve as an enabler for greater inter-team mobility. It aligns the experiment-driven consumer-facing product development workflow with how internal Infrastructure teams operate, thus making it easier to collaborate or switch between teams.
@@ 9.4. Avoiding metric-blindness
At the same time it is important to recognize that metric movement should not become a condition or even barrier to pursuing revolutionary ideas. Some improvements may not result in (measurable) DAT reductions, either because they cannot be measured or because no valid experiment can be setup. It remains paramount that we as software builders continue to listen to our expert opinion on future innovations as well as current workflow hindrances, next to assessing more large-scale sentiment and qualitative feedback from fellow developers. We must maintain a balance between leveraging DAT for data-driven decisions and fostering a culture where innovative ideas can thrive. Consider the example of coding agents that ingest a development task and work on it in a human-like way (Hassan et al., 2024). Due to their current speed limitations, they might initially have adverse impact on DAT (i.e., a developer can perform the same actions in a much shorter DAT-equivalent than the agent), though their long-term trajectory looks extremely promising.
Despite these challenges, DAT provides key insights into developer velocity, and is measured against when improving developer workflows. By precisely capturing the previously opaque authoring time, DAT enabled a paradigm shift for Infrastructure developers, bringing them into a more data-driven and objective world.
9.5. Experimenting at Industry-Scale
While a relatively large body of such studies exists in the space of types (Hanenberg, 2010; Endrikat et al., 2014; Kleinschmager et al., 2012), understandably, one limitation of many previous studies was the reliance on students as experiment subjects (Carver et al., 2004) and the inability to do like-for-like comparisons between control and experiment groups. This might be a reason why studies came to contradictory results on the effectiveness of types. Our experiment here suffers no such limitations and is, to the best of our knowledge, the first one to be conducted in a large-scale, state-of-the-art, industrial setting with real developers on real-world problems. Partially contradicting previous findings (Hanenberg, 2010), our results show strong productivity improvements when using types. It is possible that the nature of our experiment, namely the introduction of typed mocking in a large-scale system, is more comparable to previous studies which favored types as system or task complexity increased (Stuchlik and Hanenberg, 2011; Endrikat et al., 2014; Gao et al., 2017). While this hypothesis seems very plausible, we need more rigorous experiments in different settings to assess whether static types hinder ad-hoc, quick development, but are actually a productivity improvement in larger projects. It is also unclear where the threshold between these two extremes lies.
For code sharing and compiler improvements, there exists as body of largely anecdotal or post-hoc empirical analyses touting their benefits: These range from cost reduction to (Haefliger et al., 2008; Bhattacharya and Neamtiu, 2011; Mockus, 2007). However, these works are often somewhat older, seldom contain concrete efficiency gains, stem mostly from Open-Source Systems, or lack a focus on (time-based) developer productivity. Moreover, we are unaware of controlled experiments, making the studies in this paper possibly the first in the respective areas.
10. Future Work
As AI advances in software development (Khemka and Houck, 2024), we believe DAT will remain relevant. Coding agents learn from human interactions, and DAT could guide the optimization (or loss) function toward higher productivity.
11. Conclusion
@@
    Increased use of code sharing, showing a $>50\%$ DAT improvement, saving thousands of hours annually.
In conclusion, DAT has enabled Meta’s Infrastructure organization to adopt a metric-first development approach, aligning it closer with product teams. This way, DAT enabled a paradigm shift in how Infrastructure engineers work, report, and plan. DAT also paved the way toward rigorous experimentation on long-standing software engineering debates, such as “do types make development more efficient,” marking significant business and research advancements.

### v2

@@ 6.3. Results
We initially applied this method to two major code-sharing frameworks, covering thousands of code-sharing diffs and counterfactual diffs monthly (total $n=\sim 50,000$). Results indicate that for these two frameworks alone, we save thousands of hours of DAT annually, a relative improvement to non-shared development of over 50%. We then expanded the methodology with more parameters to other frameworks. Results generally pointed to large engineering savings from code sharing on a continuous basis.
@@ 7. Threats to Validity
7.1. Construct Validity
To mitigate implementation errors, we performed a quadruple validation of DAT (see Section 3.3), including large-scale surveys and end-to-end tests with GT. Moreover, we plan to increase accuracy by supporting more tools and IDEs with precise matches. For instance, we have already seen the precision improvements of the addition of Android Studio as a natively supported IDE.
7.2. Internal Validity
As a top-line metric, DAT strikes a balance between support for all engineering workflows at the company (recall), and correctness for each covered diff (precision). These objectives are often at odds, as the example of Anchor-DAT shows: extrapolation increases tool coverage at the expense of potentially covering activity not related to this diff.
To cater to these different demands, we designed a highly accurate base algorithm for DAT (precise matches), and augment it with a heuristic. DAT consumers can choose their individual precision-recall trade off based on their requirements. DAT integrates with Meta’s internal experimentation framework Deltoid, which automatically handles power analysis, effect size calculation, experiment tracking, multiple simultaneous experiment adjustments, and reporting (Facebook Engineering, 2014), ensuring reported effects uphold high rigor.
7.3. External Validity
We believe DAT can apply to a multitude of stakeholders outside of Big Tech. It might be easier to obtain buy-in in other organizations than Meta, which needs to support a plethora of products, environments, and developer preferences. For example, an organization that develops a Windows app would only need to support the exact tools used in their specific development flow.
While experimentation is most effective with enough data points in the control and test group, it is often easier to run case studies, which can also provide convincing evidence. Power analysis can reveal that sometimes a few dozen developers can be sufficient to detect effect sizes of the magnitude we present in this paper (Fitzner and Heckinger, 2010). It is also not strictly necessary to have full control over the complete engineering workflow, as one is typically interested in a specific aspect of it with an experiment. There are many organizations that have the means to acquire such data and drive improvements based on it, such as consultancies, tool vendors, and open source projects (for example, the Eclipse IDE for Java surpassed 12 million monthly downloads in 2023 (Krill, 2023)). Even some researchers (Beller et al., 2017) attract developer bases large enough to run sufficiently powered experiments. Finally, even in the absence of many users, can the mental model of experimentation-guided software development help shape how developers build products.
@@ 8. Discussion
8.1. Studies at Industry Scale
8.1.1. Experiments
While a relatively large body of such studies exists in the space of types (Hanenberg, 2010; Endrikat et al., 2014; Kleinschmager et al., 2012), understandably, one limitation of many previous studies was the reliance on students as experiment subjects (Carver et al., 2004) and the inability to do like-for-like comparisons between control and experiment groups. This might be a reason why studies reached contradictory conclusions about the effectiveness of types. Our experiment here suffers no such limitations and is, to the best of our knowledge, the first one to be conducted in a large-scale, state-of-the-art, industrial setting with real developers on real-world problems. Partially contradicting previous findings (Hanenberg, 2010), our results show strong productivity improvements when using static types. It is possible that the nature of our experiment, namely the introduction of typed mocking in a large-scale system, is more comparable to previous studies which favored types as system or task complexity increased (Stuchlik and Hanenberg, 2011; Endrikat et al., 2014; Gao et al., 2017). While this hypothesis seems plausible, we need more rigorous experiments in different settings to assess whether static types hinder ad-hoc, quick development, and are actually a productivity improvement in larger projects. It is also unclear where the threshold between these two extremes lies.
8.1.2. Case Studies
For code sharing and compiler improvements, there exists a vast body of largely anecdotal or post-hoc empirical analyses touting their benefits: These range from cost reduction to increased developer satisfaction (Haefliger et al., 2008; Bhattacharya and Neamtiu, 2011; Mockus, 2007). These works corroborate our findings in spirit, but seldom contain concrete efficiency gains, stem mostly from Open-Source Systems, lack a focus on (time-based) developer productivity, or are limited in size. The studies in this paper address these gaps by establishing concrete productivity gains, working on industrial systems, and containing a large amount of data.
8.2. Practical Implications
Practitioners in platform and framework teams often ask how they can affect developer productiveness and satisfaction. The studies in this paper show that these teams can have significant impact on the productivity of the engineers in their organization, assuming they focus their efforts on the right projects. By contrast, other studies performed with DAT and not reported in this paper showed that small, incremental UX improvements often did not significantly affect DAT.
This distinguishes internal developer experimentation from external consumer experimentation, where changing the position of a button by a few pixels can yield measurable changes in user behavior (Goodui.org, 2020). The fidelity of DAT might not lend itself to such micro optimizations.
This could be attributed to the fact that developers are highly cognitively engaged in their work and constantly striving to find optimizations themselves. Their scope and creativity are broader than the often confined ecosystems in the consumer-facing product world, which means that capturing workflows in an accurate metric is much harder. An example would be developers that continue to use their favorite editor, even if there is only one “officially” supported editor, and paste their edits in. The DAT for such copy-and-pasted diffs would be artificially low, and would introduce noise.
8.3. Dealing with Small Effect Sizes
Another challenge is establishing statistical significance in DAT, especially when working with small sample sizes. To overcome this, experiments often need to run for extended periods to account for variability in workflows and usage patterns.
In such cases, engineering teams might show impact by improving an operational metric more directly affected by their changes, for example, number of successful searches. While a statistically significant, positive change in this metric might not yield statistically significant effects on DAT, the team could rely on a previously demonstrated, principal correlation between that metric and DAT to claim productivity improvements.
8.4. Embracing an Experiment-first Culture
Setting up an experiment requires time and rigor, and may require a shift in the behavior of engineers, who might not be used to this mindset. As an example, the experiment in Section 4 involved a staged migration of mock call sites, to ensure DAT would be correctly measured. This additional effort for the sake of an experiment cannot always be justified. To design code in a DAT-first way can significantly alter the feature itself.
As a benefit this mindset aligns the experiment-driven consumer-facing product development workflow with how internal Infrastructure teams operate, thus making it easier to collaborate or switch between teams. Thus, DAT can facilitate greater inter-team mobility and more uniform workflows in the company.
@@ 8.5. Going Beyond Velocity
This paper demonstrates the broad applicability of DAT in practice, enabling Software Engineering-related case studies and experiments, from language features and compiler improvements to a generalized method to estimate framework efficiency gains. While DAT offers insights into velocity, there are other important aspects to software development such as quality (Cavano and McCall, 1978) or developer satisfaction (Storey et al., 2019). Our experiments and case studies have shown, though, that these often go together: a DAT improvement frequently co-occurs with other desired properties, such as more static safety guarantees through typed mocking in Section 4 or higher developer satisfaction.
Our studies have established that language and framework improvements (e.g., the React compiler improvements in Section 5) provide large velocity improvements. However, for an end-to-end analysis, their initial development efforts need to be balanced against the later DAT savings, ensuring the benefits outweigh the initial added development overhead. Fortunately the initial development cost can be measured with DAT, too, providing a net benefit analysis of these features. Having performed over 20 DAT experiments and case studies, we are now in a position to inform planning and prioritization of development in a rigorous, data-driven way.
@@ 8.6. Avoiding Metric-Blindness
DAT has provided key insights into developer velocity. By precisely capturing the previously opaque authoring time, DAT enabled a paradigm shift for Infrastructure developers, bringing them into a more data-driven and objective world.
Yet, it is important to recognize that metric movement should not become a condition or barrier to pursuing revolutionary ideas. Some improvements may not result in (measurable) DAT reductions, either because they cannot be measured or because no valid experiment can be set up. It remains paramount that we as software builders continue to listen to our expert opinion on future innovations as well as current workflow hindrances, next to assessing more large-scale sentiment and qualitative feedback from fellow developers. We must maintain a balance between leveraging DAT for data-driven decisions and fostering a culture where innovative ideas can thrive. Consider the example of coding agents that ingest a development task and work on it in a human-like way (Hassan et al., 2024). Due to their current speed limitations, they might initially have adverse impact on DAT (i.e., a developer can perform the same actions in a much shorter DAT-equivalent than the agent), though their long-term trajectory looks extremely promising.
9. Conclusion & Future Work
@@
    Increased use of code sharing, showing a $50\%$ DAT improvement, saving thousands of hours annually.
In conclusion, DAT has enabled Meta’s Infrastructure organization to adopt a metric-first development approach, aligning it closer with product teams. This way, DAT instigated a paradigm shift in how Infrastructure engineers work, report, and plan. DAT also paved the way toward rigorous experimentation on long-standing software engineering debates, such as “do types make development more efficient.”
As AI advances in software development (Khemka and Houck, 2024), we believe DAT will remain relevant. Coding agents learn from human interactions, and DAT could not only be used to measure continuous improvements, but also guide the optimization (or loss) function toward higher productivity. Finally, the shift to a scientific, experiment-driven internal engineering culture will hopefully have long-lasting impact on engineers and organizations.